a) What is a well-behaved wave function? Illustrate with diagrams. (5)
b) Write the values of all four quantum numbers for 2p electrons. Discuss their
significance in brief.
1
Expert's answer
2019-03-27T04:36:26-0400

a. Wave function is a complex function. It describes the state of the particle at all points of space and at any time. To understand the physical content of the wave function, we turn to experiments on electron diffraction. It turns out that clear diffraction patterns are detected even if you send single electrons to the target, that is, when each subsequent electron is released after the previous reaches the screen. After sufficient prolonged bombardment, the picture on the screen will exactly correspond to the one that comes out simultaneously sending a large number of electrons to the target. From this it can be concluded that the motion of any microparticle separately, including the location of its detection, is subject to statistical (probabilistic) laws, and when directed to the target of a single electron point on the screen, in which it will be fixed, in advance from 100% -that it is impossible to predict with certainty.


b. n=2; l=1; ml= −1,0,+1; ms= −1/2 or +1/2

Principle Quantum Numbers: This quantum number describes the size of the orbital. It is represented by n. 

 Azimuthal Quantum Number: This quantum number describes the shape of the orbital. It is represented as 'l'. The value of l ranges from 0 to (n-1). For l = 0,1,2,3... the orbitals are s, p, d, f..

Magnetic Quantum Number: This quantum number describes the orientation of the orbitals. It is represented as ml. The value of this quantum number ranges from . When l = 2, the value of ml will be -2, -1, 0, +1, +2.

  Spin Quantum number: This quantum number describes the direction of electron spin. This is represented as ms. The value of this is +1/2 for upward spin and -1/2 for downward spin.